A rectangular field, one of whose sides is the straight edge of a river is to be enclosed by 600 metres of fencing on the remaining three sides. what would be the length and breadth of the rectangle if encloased area is to be as large as possible.
 
 
please answer the above problem as soon as possible.

Grade:11

1 Answers

Susmita
5 years ago
All right.So if l is length,b is breadth(or vice versa) then according to the problem
2l+b=600.
Or,b=600-2l
Area A=bl=(600-2l)l=600l-2l2
For maxima or minima
dA/dl=0
Or,600-4l=0
Or,l=150
So b=300
For maxima d2A/dl2
Now d2A/dl2=-4
please approve if helped.